Institut for Matematik og Datalogi Odense Universitet March 26, 1996

JFB

DM11 – F96 – Lecture 8

Lecture, March 20

We covered sections 4.4 and 4.5 of chapter 4, plus some more on probability.

Lecture, March 27

We will cover sections 4.6, 4.7, 5.4 and 5.5.

Lecture, April 10

We will cover sections 5.1 and 5.2.

Problems to be discussed in week 14

• 15b, 18, 20, and 36 from section 4.6 .

• 14 from section 4.7.

• 6, 10, 12, 18, 24, 26 from section 5.4.